OpenStudy (anonymous):

one side of a rectangle is 5 inches and the other side is x inches. find the value of x if the area must be at least 60 square inches

OpenStudy (anonymous):

\[Area=5\times{x}=60\]

OpenStudy (anonymous):

A=l*x 60inch=5inch*x x=60/5 x=12inches

OpenStudy (anonymous):

\[x=12\]

OpenStudy (anonymous):

area of rectangle is L*W. substituting the given values. i came up with 60 = 5*x. x= 60/5= 12. therefore, x is greater than or equal to 12.